# Who’s in and who’s out? A case study of multimodal CLIP-filtering in DataComp

### v1

@@ 7.2. News sites
Figure 11. New sites with the highest pass rates (top) and lowest pass rates (bottom). News platforms that correspond to non-Western regions are highlighted in yellow, and we see high presence among news sites with lowest pass rates.
@@ 7.4. Origin date
The Wayback Machine API (Archive, 2022) responses indicate that the Internet Archive does not record snapshots for a majority of the exact URL addresses in our subsample — we only track index dates for 22.7% of the one million CommonPool samples. Of those tracked, we find that these samples overwhelmingly come from sources that were first indexed within the last ten years. We also find that pass rates tend to be somewhat higher the more recent the year. We defer the graphs and details to Section B.4.
@@ 8. Discussion
Our empirical analysis demonstrates that CLIP-filtering on the DataComp CommonPool dataset removes data nonuniformly when broken down by demographic group imputation, geographic region, or source domain. First, we find that the CLIP filter disproportionately excludes data relating to various proxies for demographic groups which include LGBTQ+ identities, women, and non-Western regions. Second, these filter discrepancies often amplify representation biases present in the raw dataset, where underrepresented groups are filtered out at higher rates compared to overrepresented groups. Third, we find that data from sites where the images themselves are the intellectual property (e.g., stock photos and educational curriculum) are considered high-quality by CLIP and kept at higher rates. Finally, the presence of NSFW samples after an initial cursory search of a small subset of CommonPool reveals that the combination of applying NSFW and CLIP filters fails to catch all instances of sexually-explicit content.
In this section, we elaborate on how the above findings imply that CLIP-filtering and similar filtering practices may result in societal harms and potential legal implications. Using the CLIP filter as a case study, we then dive into several key assumptions and failures of existing data cleaning methods when they are used to build large image-text models. Doing so allows us to form recommendations for designing, implementing, and evaluating data filtering techniques and data curation methods.
@@ 8.1. Societal consequences of data filtering
We highlight three aspects of CLIP-filtering CommonPool data that have societal ramifications: the amplification of demographic bias, the upweighting of websites that host licensed or copyrighted images, and the inclusion of unmoderated inappropriate content. Altogether these dimensions illustrate an example of how existing data filtering practices have failed to mitigate certain downstream harms and, hence, that stochastic neural-network-based filtering may not be the end-all-be-all technique to capture all “impurities” in web-scraped data.
@@
As mentioned in Section 4.4, we cannot necessarily jump to conclusions that all else held equal, varying the demographic group will change the filter likelihood, and we leave this type of causal investigation (Kusner et al., 2017) for our future work. It is, for example, possible that more data relating to a certain group can come from websites with established alt-text practices, which would change filtering rates even without using a biased model like CLIP. At the same time, however, the finding of correlations between imputed demographic groups and filtering rates is important absent this causal measurement. For CommonPool, the CLIP filter modifies the original demographic distribution and thus does not result in a uniform sampling of web-scraped data. Our analysis in Section 5.1 proves that CLIP-filtering can create training datasets more unrepresentative of certain populations and containing higher proportions of stereotypes than the unfiltered datasets.
@@
In the recent lawsuit from The New York Times against OpenAI, the plaintiff argues that The New York Times articles are considered more “valuable” during the training process (Grynbaum and Mac, 2023), where “value” here refers to the amount of content contributed to training a model. OpenAI in response has claimed that these articles are just “a tiny slice of overall training data” (OpenAI, 2024). Our analysis in Section 7.2 provides additional evidence that data from certain websites, including The New York Times, are implicitly upweighted after filtering, which implies that systems trained through this kind of filtering do in fact place more value on certain types of copyrighted works, although it is unknown if OpenAI implements a similar filtering mechanism in constructing its training data.
@@ 8.2. Assumptions in data filtering practices
In our work, we demonstrate instances of stereotypes perpetuated by the CLIP filter when examining common words or intersections of demographic dimensions (Sections 5.1.2 and 5.1.3), as well as instances of Western bias even when controlling for English text (Sections 6 and B.2). These trends support the notion that the OpenAI CLIP model is not intended to assess image-text alignment, as pointed out in prior work on image captioning CLIP evaluations (Qiu et al., 2023) and on CLIP classification bias (Agarwal et al., 2021). Especially because OpenAI CLIP was never designed to be a data filter, much less a filter used to build deployed models (OpenAI, 2022), we strongly caution against using CLIP as a filter to build future training datasets.
It is entirely plausible that OpenAI CLIP is trained on a distribution of data similar to the filtered dataset makeup, which means the CLIP filter may implicitly attempt to replicate its own training dataset. Our findings on exclusion amplification confirm this notion, as well-represented classes of data seem to be included at higher rates. However, none of this can be confirmed given that the training data for OpenAI CLIP has not been released, nor have the data filtering steps to obtain said training data been formally described (Radford et al., 2021). In other words, the CLIP model is generated by a black-box pipeline, trained on an unknown dataset. This results in a black-box filtering model whose behavior is difficult to predict without prohibitively many interactions with it. The main advantage of CLIP-filtering here is scale, which embodies the notion of how “scale beats noise” presented in Birhane et al. (2021).
@@ 8.3. Recommendations
After questioning the assumptions of current data filtering practices, we subsequently form several recommendations — while most of these recommendations apply to data filtering methods specifically, we argue that that data filters cannot be treated as a panacea (Birhane et al., 2021). As such, we extend some of our recommendations to the data curation process more generally.
@@ 9. Conclusion
In this work, we audit the CLIP-filtering step in the DataComp and LAION pipelines, where to construct “high-quality” large-scale training datasets from web-scraped data, an existing pre-trained CLIP model is used to determine image-text alignment. We find that current forms of image-text filtering, similar to prior work (Gururangan et al., 2022; Lucy et al., 2024; Dodge et al., 2021), embed societal judgements in determining what kinds of data should be discarded. Specifically, we find that content relating to marginalized identities or non-Western regions are more likely to be left out of the final training dataset. Moreover, we demonstrate examples of exclusion amplification – data from certain imputed demographic groups already underrepresented in the unfiltered dataset are excluded at higher rates from the downstream dataset. We find that copyrighted data from certain types of websites are judged to be higher quality and that NSFW filters fail to remove large quantities of sexually-explicit text, raising CSAM and NCII concerns. These findings raise new issues on the downstream societal effects of filtering methods within the broader machine learning pipeline, especially as large multimodal models become increasingly deployed in high-stakes settings (Ali et al., 2023; Booth et al., 2021).
@@ A.2. CLIP kNN
    The distribution of the nearest neighbors become the probability scores of the gender or race of the DataComp image, and the argmax is the final group annotation.
@@ C.1. Remove images from websites labeled as NSFW by Cloudflare
In Section 7.3.3, our audit finds that $1.03\%$ of CommonPool samples come from websites that Cloudflare labeled as NSFW (including pornography). These samples could easily be detected and removed from the original dataset, without checking or downloading image content. While simple to implement, we again caution that this method would not remove all NSFW content and would not impact copies of the dataset already downloaded nor models already trained on CommonPool. The NSFW-categorized domains corresponding to the subset of $400,000$ samples are available upon request.

### v2

@@ 7.2. News sites
Figure 11. New sites with the highest pass rates (top) and lowest pass rates (bottom). News platforms that correspond to non-Western regions are highlighted in yellow, and we see high non-Western presence among news sites with lowest pass rates.
@@ 7.4. Origin date
The Wayback Machine API (Archive, 2022) responses indicate that the Internet Archive does not record snapshots for a majority of the exact URL addresses in our subsample — we only track index dates for 22.7% of the one million CommonPool samples. Of those tracked, we find that these samples overwhelmingly come from sources that were first indexed within the last ten years. We also find that pass rates tend to be somewhat higher the more recent the year. We present the graphs and details in Section B.4.
@@ 8. Discussion
Our analysis demonstrates that CLIP-filtering on the DataComp CommonPool dataset removes data nonuniformly when broken down by demographic group imputation, geographic region, or source domain. First, we find that the CLIP filter disproportionately excludes data relating to various proxies for demographic groups which include LGBTQ+ identities, women, and non-Western regions. Second, these filter discrepancies often amplify representation biases present in the raw dataset, where underrepresented groups are filtered out at higher rates compared to overrepresented groups. Third, we find that data from sites where the images themselves are the intellectual property (e.g., stock photos and educational curriculum) are considered high-quality by CLIP and kept at higher rates. Finally, the presence of NSFW samples after an initial cursory search of a small subset of CommonPool reveals that the combination of applying NSFW and CLIP filters fails to catch all instances of sexually-explicit content.
In this section, we elaborate on how the above findings imply that CLIP-filtering and similar filtering practices may result in societal harms and potential legal implications. Using the CLIP filter as a case study, we then dive into several key assumptions and failures of existing data cleaning methods when they are used to build large image-text models. For instance, we discuss how the CLIP filter may be more likely to pass samples similar to its own training data in Section 8.2.1. Finally, we use our analysis to inform recommendations for designing, implementing, and evaluating data filtering techniques and data curation methods.
@@ 8.1. Societal consequences of data filtering
We highlight three aspects of CLIP-filtering CommonPool data that have societal ramifications: the amplification of demographic bias, the inclusion of unmoderated inappropriate content, and the upweighting of websites that host licensed or copyrighted images. Altogether these dimensions illustrate an example of how existing data filtering practices have failed to mitigate certain downstream harms and, hence, that stochastic neural-network-based filtering may not be the end-all-be-all technique to capture all “impurities” in web-scraped data.
@@
As mentioned in Section 4.3, we cannot necessarily jump to conclusions that all else held equal, varying the demographic group will change the filter likelihood, and we leave this type of causal investigation (Kusner et al., 2017) for our future work. It is, for example, possible that more data relating to a certain group can come from websites with established alt-text practices, which would change filtering rates even without using a biased model like CLIP. At the same time, however, the finding of correlations between imputed demographic groups and filtering rates is important absent this causal measurement. For CommonPool, the CLIP filter modifies the original demographic distribution and thus does not result in a uniform sampling of web-scraped data. Our analysis in Section 5.1 shows that CLIP-filtering can create training datasets more unrepresentative of certain populations and containing higher proportions of stereotypes than the unfiltered datasets.
@@ 8.1.2. Continued presence of lightly or unmoderated sexually-explicit material raises CSAM and NCII concerns
There has been extensive discussion on the copyright implications of generative models trained on large-scale datasets scraped from the web. In the recent lawsuit from The New York Times against OpenAI, for instance, the plaintiff has argued that The New York Times articles are considered more “valuable” during the training process (Grynbaum and Mac, 2023), where “value” here refers to the amount of content contributed to training a model. OpenAI in response has claimed that these articles are just “a tiny slice of overall training data” (OpenAI, 2024). Our analysis in Section 7.2 provides additional evidence that data from certain websites, including The New York Times, are implicitly upweighted after filtering, which implies that systems trained through this kind of filtering do in fact place more value on certain types of copyrighted works, although it is unknown if OpenAI implements a similar filtering mechanism in constructing its training data.
In order to make a case for copyright infringement in the United States, according to the fair use doctrine, the defendant must demonstrate that use of copyrighted material follows four factors (of the United States, 1976). According to legal analysis of fair use in the context of generative machine learning models, the fourth factor of fair use, which examines the potential market effect, draws the most attention from legal scholars (Lee et al., 2023; Henderson et al., 2023). Alhadeff et al. (2024), for example, argue that the fourth factor weighs against a finding of fair use because the output generative AI tools may be a substitution for market harm.
@@ 8.2. Assumptions in data filtering practices
In our work, we demonstrate instances of stereotypes perpetuated by the CLIP filter when examining common words or intersections of demographic dimensions (Sections 5.1.2 and 5.1.3), as well as instances of Western bias (Section 6). These trends support the notion that the OpenAI CLIP model is not intended to assess image-text alignment, as pointed out in prior work on image captioning CLIP evaluations (Qiu et al., 2023) and on CLIP classification bias (Agarwal et al., 2021). Especially because OpenAI CLIP was never designed to be a data filter, much less a filter used to build deployed models (OpenAI, 2022), we strongly caution against using CLIP as a filter to build future training datasets.
It is entirely plausible that OpenAI CLIP is trained on a distribution of data similar to the filtered dataset makeup, which means the CLIP filter may implicitly attempt to replicate its own training dataset. Our findings on exclusion amplification confirm this notion, as well-represented classes of data seem to be included at higher rates. However, none of this can be confirmed given that the training data for OpenAI CLIP has not been released, nor have the data filtering steps to obtain said training data been formally described (Radford et al., 2021). In other words, the CLIP model is generated by a black-box pipeline, trained on an unknown dataset. This results in a black-box filtering model whose behavior is difficult to predict without prohibitively many interactions with it. The main advantage of CLIP-filtering here is automation, which embodies the existing notion of how “scale beats noise” as characterized in Birhane et al. (2021).
@@ 8.3. Recommendations
After questioning the assumptions of current data filtering practices, we subsequently form several recommendations. While most of these recommendations apply to data filtering methods specifically, we argue that that data filters cannot be treated as a panacea (Birhane et al., 2021). As such, we extend some of our recommendations to the data curation process more generally.
@@ 9. Conclusion
In this work, we audit the CLIP-filtering step in the DataComp and LAION pipelines, where to construct “high-quality” large-scale training datasets from web-scraped data, an existing pre-trained CLIP model is used to determine image-text alignment. We find that current forms of image-text filtering, similar to prior work (Gururangan et al., 2022; Lucy et al., 2024; Dodge et al., 2021), embed societal judgements in determining what kinds of data should be discarded. Specifically, we find that content relating to marginalized identities or non-Western regions are more likely to be left out of the final training dataset. Moreover, we demonstrate examples of exclusion amplification – data from certain imputed demographic groups already underrepresented in the unfiltered dataset are filtered out at higher rates. We find that copyrighted data from certain types of websites are judged to be higher quality and that NSFW filters fail to remove large quantities of sexually-explicit text, raising CSAM and NCII concerns. These findings raise new issues on the downstream societal effects of filtering methods within the broader machine learning pipeline, especially as large multimodal models become increasingly deployed in high-stakes settings (Ali et al., 2023; Booth et al., 2021).
@@ A.2. CLIP kNN
    The distribution of the nearest neighbors becomes the probability scores of the gender or race of the DataComp image, and the argmax is the final group annotation.
@@ C.1. Remove images from websites labeled as NSFW by Cloudflare
In Section 7.3.3, our audit finds that $1.03\%$ of CommonPool samples come from websites that Cloudflare labeled as NSFW (including pornography). These samples could easily be detected and removed from the original dataset, without checking or downloading image content. While simple to implement, we again caution that this method would not remove all NSFW content and would not impact copies of the dataset already downloaded nor models already trained on portions of CommonPool. The NSFW-categorized domains corresponding to the subset of $400,000$ samples are available upon request.